OpenStudy (anonymous):

Can someone tell me if I solved this trinomial correctly? 2x^2 + 13x + 15. The first trinomial with be 2x^2 + 13x + 15. To solve you first you check to see if there is a GCF. For this one there is none. Next you are going to multiply the first coefficient which is 2 with the last term of the trinomial which is 15. So 2 * 15 = 30. Next you need to find factors of 30 that add to give you the coefficient of the middle term which is 13. I used -2 + 15. So now it is, 2x^2 - 2x + 15x + 15. Now I need to separate it into two groups giving me, 2x^2 - 2x and 15x + 15. Now I need to find the GCF of

OpenStudy (anonymous):

the first group which is 2x. And the GCF of the second group is 15. So now it is 2x(x - 1) 15(x + 1). Finally I factor the common binomial which is (x - 1)(2x + 15).

OpenStudy (vincent-lyon.fr):

Use the "Mathematics" section, not "Language and Culture" for this question. You will get an answer much sooner.

OpenStudy (lgbasallote):

not to mention more accurate and detailed

OpenStudy (anonymous):

Did your teacher ever show you the FOIL method? FOIL is an acronym. FIRST OUTSIDE INSIDE LAST FOIL shows you the variables, from each set, that you need to multiply. I'll show you how to use FOIL.... ~~~~~~~ Look at your answer: (x - 1) (2x + 15) 1. Multiply the FIRST variables from each set......x * 2x = 2x^2 2. Multiply the OUTSIDE variables from each set.....x * 15 = 15x 3. Multiply the INSIDE variables from each set.....-1 * 2x = -2x 4. Multiply the LAST variables from each set.....-1 * 15 = -15 Now....combine all your answers....and you get.... 2x^2 + 15x -2x - 15.....which is the same as.... 2x^2 + 13x -15. BUT THE TRINOMIAL YOU WANT is 2x^2 + 13x + 15 So you see how your answer isn't quite right? It's very close though. ~~~~~~~~~ The answer that I came up with is....(2x + 3) (x + 5) And if you use the FOIL method for those two groups, you'll see it the trinomial is 2x^2 + 13x + 15, which matches the trinomial you want.